• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name riordan

    Log entry time 17:33:45 on October 04, 2010

    Entry number 333631

    keyword=Run Plan for Swing 10/4 through Day 10/5

    Jack and company started looking at the magnets today and will need a few more days to get everything ready. Dave Meekins is working on the target system to get it ready for a cooldown tomorrow.

    Plans for the next 24 hours:

    JP will cool down the target tomorrow. We intend to have LH2 and LD2. If the dipole becomes available tomorrow we will see if we can start HRS commissioning. Compton tuning has been delayed (needs to be rescheduled, I'm shooting for Tuesday). BCM calibrations were delayed until 07:00 Wednesday

    We determined today that the hall's energy will drop by about 10 MeV when Hall C comes up, so any further Moller measurements or beam energy measurements will have to be redone. Arc measurement is postponed until then.

    I'll send out an announcement to the collaboration regarding target shift status once the target cooldown starts.

    Work to be scheduled:

    A copy of this log entry has been emailed to: riordan, hyde, camsonne, jroche, munoz